how long would it take you to walk 10km if your time was 6.6 seconds in 12 meters?
Answer by mananth(16946)

6.6 seconds 12 m
1 sec= 12/6.6 = 1.82m/s
...
10km = 10,000 m
speed = 1.82m/s
time = 10000/1.82
=5494 seonds
=5494/3600
=1.53 hours